OpenStudy (anonymous):

x ^{sqrt{logx}}=10^{8}

OpenStudy (anonymous):

\[x ^{\sqrt{logx}}=10^{8}\]

OpenStudy (anonymous):

how does 8log10 = 8?

OpenStudy (anonymous):

what are we doing here? just solving for x?

OpenStudy (anonymous):

yeah without using a calculator

OpenStudy (anonymous):

that is because log (10) = 1, when there is no base, its understood to be base 10

OpenStudy (anonymous):

but ten wasn't the base

OpenStudy (anonymous):

then what is the base? you gotta put it or else people will think its 10

OpenStudy (anonymous):

ok sorry let me show you where i got to

OpenStudy (anonymous):

\[\left( \sqrt{logx} \right)\left( logx \right)=10^{8}\]

OpenStudy (anonymous):

\[=8\log10\]

OpenStudy (anonymous):

Right. So you are taking the log (base 10) of both sides of the equation. the right hand side simplifies to 8

OpenStudy (anonymous):

So you should have: \[\sqrt{\log(x)}*\log(x) = 8\]

OpenStudy (anonymous):

does it simplify because since the logs base is ten and since its \[8log _{10}10\] thelog simplifies to 1?

OpenStudy (anonymous):

Right :)

OpenStudy (anonymous):

ohhhh thanks see i knew the base of log was ten but i didnt know that i could do that thanks!!

OpenStudy (anonymous):

So do you have an idea of where to go from here?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

cool, have fun :)